# TradExpert: Revolutionizing Trading with Mixture of Expert LLMs

### v1

@@ Abstract
The integration of Artificial Intelligence (AI) in the financial domain has opened new avenues for quantitative trading, particularly through the use of Large Language Models (LLMs). However, the challenge of effectively synthesizing insights from diverse data sources and integrating both structured and unstructured data persists. This paper presents TradeExpert, a novel framework that employs a mix of experts (MoE) approach, using four specialized LLMs, each analyzing distinct sources of financial data, including news articles, market data, alpha factors, and fundamental data. The insights of these expert LLMs are further synthesized by a General Expert LLM to make a final prediction or decision. With specific prompts, TradeExpert can be switched between the prediction mode and the ranking mode for stock movement prediction and quantitative stock trading, respectively. In addition to existing benchmarks, we also release a large-scale financial dataset to comprehensively evaluate TradeExpert’s effectiveness. Our experimental results demonstrate TradeExpert’s superior performance across all trading scenarios.
Introduction
@@
Before the advent of LLMs, traditional financial models (Zeng et al. 2023; Yang et al. 2020; Liu et al. 2020; Baek and Kim 2018), primarily relied on quantitative methods such as statistical analysis, time series forecasting, and econometric models. These models often struggled to incorporate unstructured data such as news articles or financial reports without manual intervention. As a result, the development of LLMs tailored for financial applications has progressed rapidly. Initial ventures into this domain repurposed general LLMs such as GPTs (Brown 2020; Achiam et al. 2023) and LLaMAs (Touvron et al. 2023a, b; AI@Meta 2024) to interpret financial texts. However, more specialized language models such as FinBERT (Araci 2019), BloombergGPT (Wu et al. 2023), and FinGPT (Yang, Liu, and Wang 2023) have since evolved, demonstrating enhanced proficiency in understanding and predicting market movements from unstructured data. These models were specifically fine-tuned or pre-trained on vast amounts of financial corpus. This extensive training on domain-specific datasets has allowed them to better capture typical patterns in the financial corpus. Despite these advancements, the challenge remains to effectively synthesize insights from diverse data sources like historical stock prices, alpha factors, fundamental data, news articles, etc. In addition, integration of the deluge of unstructured financial texts with structured quantitative metrics still remains to be investigated with language models.
@@
To this end, we propose the TradExpert framework, which stands at the confluence of these challenges. It leverages a Mixture of Experts (MoE) approach (Eigen, Ranzato, and Sutskever 2013; Du et al. 2022; Shen et al. 2023), involving multiple LLMs each specialized in distinct facets of financial data—news articles, market data, alpha factors, and fundamental data. This not only enhances the model’s ability to process diverse data modalities but also allows for a more nuanced understanding of how different factors interact to influence market trends. Figure 1 illustrates differences among traditional, LLM-based, and MoE LLMs-based financial models. In TradeExpet, each expert works with a distinct focus and produces specialized reports, which are finally summarized and analyzed by a general expert, just like the structured division of labor seen in the real world. Specifically, TradExpert employs specialized LLMs to first independently analyze different data sources, then integrates these analyses via another LLM that synthesizes insights to predict market movements and inform trading strategies. Innovatively, we utilize a reprogramming mechanism to convert time series data to embeddings aligned with LLMs. In addition, we propose two modes for the General Expert LLM, prediction mode and ranking mode, for stock movement prediction and stock trading strategies, respectively. In ranking mode, we innovatively let the LLM serve as a comparator within a relaxed sorting algorithm, enabling the selection of the Top-K ranked stocks for trading. To comprehensively evaluate our method, we have also collected a large-scale datasets, which will be publicly released.
@@
have significantly advanced in recent years, blending NLP techniques with financial analytics to extract meaningful insights from vast amounts of unstructured financial data. To begin with, FinBert (Araci 2019) is a financial domain-specific variant of BERT, pretrained on a large corpus of financial communications. In 2023, BloombergGPT (Wu et al. 2023) emerged as a 50-billion-parameter model trained on a vast financial dataset. FLANG (Shah et al. 2022) introduced a financial language model with specialized masking and objectives. Astock (Zou et al. 2022) provided a platform for studying NLP-aided stock auto-trading algorithms on the Chinese market. BBT-FinT5 (Lu et al. 2023) advanced Chinese financial NLP with a large-scale pre-trained model. FinMA (Xie et al. 2023) showcased a model fine-tuned on a multi-task instruction datasets. FinGPT (Yang, Liu, and Wang 2023) provided an open-source framework for financial LLMs. InvestLM (Yang, Tang, and Tam 2023) showed the effectiveness of instruction tuning for investment-related tasks. FinReport (Li et al. 2024b) introduced a system for automatic financial report generation. Lastly, AlphaFin (Li et al. 2024a) integrated retrieval-augmented generation techniques for financial analysis. Collectively, these works demonstrate the evolution of financial NLP models and benchmarks, advancing the capabilities of LLMs in financial applications.
@@
has also been rapidly developed for stock movement prediction. StockNet (Xu and Cohen 2018) developed a deep generative model that jointly exploits text and price signals for stock movement prediction. SLOT (Soun et al. 2022) improved upon this by using self-supervised learning to handle sparse and noisy tweet data, capturing multi-level price trends. CH-RNN (Wu et al. 2018) introduced a hybrid deep sequential modeling approach that leverages social text for stock prediction, incorporating cross-modal attention mechanisms. More recently, studies (Lopez-Lira and Tang 2023; Chen et al. 2023) have explored the use of ChatGPT for stock movement prediction, comparing its performance with traditional state-of-the-art models. These works collectively demonstrate the increasing sophistication of models that integrate text and financial data, highlighting the potential for improving trading scenarios.
@@
Stastics
News
Market Data
Alpha Factors
Fundamental Data
Data Split
@@
Methodology
@@
In TradExpert, all expert LLMs are built on the LLaMA-2-7B backbone LLM (Touvron et al. 2023b) and are supervised and fine-tuned using the LoRA mechanism (Hu et al. 2022). Before training and fine-tuning, we preprocess the raw datasets to construct prompts, instructions, and ground-truth responses for each LLM. An overall description of the preprocessed datasets is demonstrated in Table 1. The details will be introduced in the following.
News Analyst
The News Analyst LLM is designed to analyze texts of news articles to predict stock movements. The prompt and instruction for fine-tuning the LLM are shown in Figure 3. The outputs from the News Analyst LLM include not only a prediction of the stock movement but also a reasoning of how the news article relates to the predicted movement in order to employ a Chain-of-Thought (CoT) (Wei et al. 2022) reasoning approach. The ground-truth reasonings are pre-generated by the OpenAI GPT-4 API using instructions and prompts that incorporate the actual stock movements and the texts of news articles.
@@
Market Analyst
The Market Analyst LLM focuses on analyzing historical OHLCV (Open, High, Low, Close, Volume) data to predict stock movements. However, time series data is inherently continuous and lacks the discrete token structure that LLMs are designed to process. This misalignment poses a significant challenge in effectively utilizing LLMs on time series. To this end, we utilize a reprogramming mechanism (Jin et al. 2024) to reprogram the input financial time series into text prototype representations.
Formally, let an OHLCV data instance be $\mathbf{X}^{(i)}\in\mathbb{R}^{N\times T}$ which consists of $N$ variables across $T$ time steps. $\mathbf{X}^{(i)}$ is first divided and embedded into a sequence of patch embeddings $\left\{\mathbf{X}_{P}^{(i)}\in\mathbb{R}^{N\times L_{P}\times d_{m}}\right\}$, where $L_{P}$ and $d_{m}$ are the number of patches and the patch embedding dimension respectively. The patches are then reprogrammed using a collection of text prototypes $\mathbf{E}^{\prime}\in\mathbb{R}^{V^{\prime}\times D}$, which is achieved by linearly probing the LLM’s pre-trained word embedding $\mathbf{E}\in\mathbb{R}^{V\times D}$, where $V$ and $V^{\prime}$ are the size of the vocabulary of the LLM and the text prototypes ( $V^{\prime}\ll V$), and $D$ is the embedding dimension. The reprogrammed patches are generated using a multi-head cross-attention mechanism: $\mathbf{Z}_{k}^{(i)}=\text{Softmax}\left(\frac{\mathbf{Q}_{k}^{(i)}\mathbf{K}_{k}^{(i)\top}}{\sqrt{d_{k}}}\right)\mathbf{V}_{k}^{(i)}$, where query $\mathbf{Q}_{k}^{(i)}=\mathbf{X}_{P}^{(i)}\mathbf{W}_{k}^{Q}$, key $\mathbf{K}_{k}^{(i)}=\mathbf{E}^{\prime}\mathbf{W}_{k}^{K}$, and value $\mathbf{V}_{k}^{(i)}=\mathbf{E}^{\prime}\mathbf{W}_{k}^{V}$ for each head $k$. The reprogrammed embeddings $\mathbf{O}^{(i)}$ are obtained by aggregating the outputs from each attention head and projecting them to the hidden dimensions of the backbone LLM. Finally, the reprogrammed embeddings are augmented with a language description of statistics extracted from TSFresh (Christ et al. 2018), serving as prompts for the Alpha Expert. An example of instruction and prompt is shown in Figure 4.
@@
For each stock, we first calculate the values of all alpha factors based on OHLCV data and then derive a comprehensive score via a LightGBM-based model (Ke et al. 2017). Subsequently, we select Top-K alphas that contribute most significantly to this comprehensive score. Descriptions of these Top-K alphas are retrieved from the database and, along with the calculated values, are included in the prompts and instructions for the Alpha Expert.
Fundamental Analyst
@@
In ranking mode, used for stock trading, the General Expert LLM functions as a comparator to establish the ranking ability. Specifically, given the summarized reports of two stocks, the General Expert LLM would determine which stock is likely to perform better in the future. To generate a Top-K ranking of stocks, we employ a relaxed comparison-based sorting similar to BubbleSort: We initially compare every pair of stocks and count the number of wins for each stock. Subsequently, we sort these counts to establish the rankings for stocks. Although algorithms like QuickSort and vanilla BubbleSort offer fewer comparisons for Top-K selection on average $\mathcal{O}(N\log N)$ and $\mathcal{O}(N\cdot K)$, we propose to use this relaxed comparison-based sorting alogrithm with $\mathcal{O}(N^{2})$ due to the non-transitive nature of LLM-based comparator (Liu et al. 2024). Therefore, more comparisons tend to yield more accurate rankings in practice.
@@
  • Benchmark Datasets: We use publicly available benchmark datasets in stock movement prediction research including CIKM18 (Wu et al. 2018), ACL18 (Xu and Cohen 2018), and BigData22 (Soun et al. 2022) datasets.
@@
Experimental Setup
In our experiments, the four expert LLMs and the General Expert LLM are bulit on the LLaMA-2-7B bakcbone model (Touvron et al. 2023b) and are finetuned via LoRA (Hu et al. 2022) mechanism.
@@
Baselines
For stock movement prediction, the baselines include: (1) Hybrid Models: StockNet (Xu and Cohen 2018), ALSTM-W (Soun et al. 2022), ALSTM-D (Soun et al. 2022), SLOT (Soun et al. 2022). 2) Large Language Models: GPT-4 (Achiam et al. 2023), Gemini (Team et al. 2023), LLaMA2-70B (Touvron et al. 2023b), LLaMA3-8B (AI@Meta 2024), FinMA-7B (Xie et al. 2023), FinGPT-LlaMA2-7B (Yang, Liu, and Wang 2023), InternLM-7B (Cai et al. 2024), Falcon-7B (Almazrouei et al. 2023), Mixtral-7B (Jiang et al. 2023).
For stock trading simulation, the baselines include: (1)Traditional Models: Random Forest (Breiman 2001), Decision Tree (Loh 2011), SVM (Cortes and Vapnik 1995). (2) Deep Learning Models: A2C (Mnih et al. 2016), PPO (Schulman et al. 2017), SARL (Ye et al. 2020), EIIE (Jiang, Xu, and Liang 2017), and DeepTrader (Wang et al. 2021). To reduce computational costs in backtesting, we evaluated all methods on datasets with stocks on the DOW 30 list, a subset of the S&P 500, with around 30 stocks.
Results
Stock Movement Prediction
Stock Trading Simulation
@@ Results
Ablation Study
The Impacts of Experts
The Effectiveness of Structured Data Reasoning.
The Choices of Ranking algorithm
@@
Conclusion
In this study, we introduced TradeExpert, a novel framework that harnesses the power of LLMs to enhance stock trading strategies. By integrating multiple specialized LLMs, each focused on distinct aspects of financial data, TradeExpert provides a comprehensive and nuanced analysis that significantly outperforms traditional financial models in practice. Looking ahead, our goal is to explore how to employ TradeExpert in the high-frequency trading scenario and extend its capabilities to encompass a wider range of global markets.

### v2

@@ Abstract
The integration of Artificial Intelligence (AI) in the financial domain has opened new avenues for quantitative trading, particularly through the use of Large Language Models (LLMs). However, the challenge of effectively synthesizing insights from diverse data sources and integrating structured and unstructured data persists. This paper presents TradExpert, a novel framework that employs a mix of experts (MoE) approach, using four specialized LLMs, each analyzing distinct sources of financial data, including news articles, market data, alpha factors, and fundamental data. The insights of these expert LLMs are further synthesized by a General Expert LLM to make a final prediction or decision. With specific prompts, TradExpert can be switched between the prediction mode and the ranking mode for stock movement prediction and quantitative stock trading, respectively. In addition to existing benchmarks, we also release a large-scale financial dataset to comprehensively evaluate TradExpert’s effectiveness. Our experimental results demonstrate TradExpert’s superior performance across all trading scenarios.
1 Introduction
@@
Before the advent of LLMs, traditional financial models Zeng et al. (2023); Yang et al. (2020); Liu et al. (2020); Baek & Kim (2018), primarily relied on quantitative methods such as statistical analysis, time series forecasting, and econometric models. These models often struggled to incorporate unstructured data such as news articles or financial reports without manual intervention. As a result, the development of LLMs tailored for financial applications has progressed rapidly. Initial ventures into this domain repurposed general LLMs such as GPTs Brown (2020); Achiam et al. (2023) and LLaMAs Touvron et al. (2023a; b); Dubey et al. (2024) to interpret financial texts. However, more specialized language models such as FinBERT Araci (2019), BloombergGPT Wu et al. (2023), and FinGPT Yang et al. (2023a) have since evolved, demonstrating enhanced proficiency in understanding and predicting market movements from unstructured data. These models were specifically fine-tuned or pre-trained on vast amounts of financial corpus. This extensive training on domain-specific datasets has allowed them to better capture typical patterns in the financial corpus. Despite these advancements, the challenge remains to effectively synthesize insights from diverse data sources like historical stock prices, alpha factors, fundamental data, news articles, etc. In addition, integration of the deluge of unstructured financial texts with structured quantitative metrics still remains to be investigated with language models.
@@
To this end, we propose the TradExpert framework, which stands at the confluence of these challenges. It leverages a Mixture of Experts (MoE) approach Eigen et al. (2013); Du et al. (2022); Shen et al. (2023), involving multiple LLMs each specialized in distinct facets of financial data—news articles, market data, alpha factors, and fundamental data. This not only enhances the model’s ability to process diverse data modalities but also allows for a more nuanced understanding of how different factors interact to influence market trends. Figure 1 illustrates differences among traditional, LLM-based, and MoE LLMs-based financial models. In TradExpert, each expert works with a distinct focus and produces specialized reports, which are finally summarized and analyzed by a general expert, just like the structured division of labor seen in the real world. Specifically, TradExpert employs specialized LLMs to first independently analyze different data sources, then integrates these analyses via another LLM that synthesizes insights to predict market movements and inform trading strategies. Innovatively, we utilize a reprogramming mechanism to convert time series data to embeddings aligned with LLMs. In addition, we propose two modes for the General Expert LLM, prediction mode and ranking mode, for stock movement prediction and stock trading strategies, respectively. In ranking mode, we innovatively let the LLM serve as a comparator within a relaxed sorting algorithm, enabling the selection of the Top-K ranked stocks for trading. To comprehensively evaluate our method, we have also collected a large-scale datasets, which will be publicly released.
@@ 2 Related Work
have significantly advanced in recent years, blending NLP techniques with financial analytics to extract meaningful insights from vast amounts of unstructured financial data. To begin with, FinBert Araci (2019) is a financial domain-specific variant of BERT, pretrained on a large corpus of financial communications. In 2023, BloombergGPT Wu et al. (2023) emerged as a 50-billion-parameter model trained on a vast financial dataset. FLANG Shah et al. (2022) introduced a financial language model with specialized masking and objectives. Astock Zou et al. (2022) provided a platform for studying NLP-aided stock auto-trading algorithms on the Chinese market. BBT-FinT5 Lu et al. (2023) advanced Chinese financial NLP with a large-scale pre-trained model. FinMA Xie et al. (2023) showcased a model fine-tuned on a multi-task instruction datasets. FinGPT Yang et al. (2023a) provided an open-source framework for financial LLMs. InvestLM Yang et al. (2023b) showed the effectiveness of instruction tuning for investment-related tasks. FinReport Li et al. (2024b) introduced a system for automatic financial report generation. Lastly, AlphaFin Li et al. (2024a) integrated retrieval-augmented generation techniques for financial analysis. Collectively, these works demonstrate the evolution of financial NLP models and benchmarks, advancing the capabilities of LLMs in financial applications.
@@
has also been rapidly developed for stock movement prediction. StockNet Xu & Cohen (2018) developed a deep generative model that jointly exploits text and price signals for stock movement prediction. SLOT Soun et al. (2022) improved upon this by using self-supervised learning to handle sparse and noisy tweet data, capturing multi-level price trends. CH-RNN Wu et al. (2018) introduced a hybrid deep sequential modeling approach that leverages social text for stock prediction, incorporating cross-modal attention mechanisms. More recently, studies Lopez-Lira & Tang (2023); Chen et al. (2023) have explored the use of ChatGPT for stock movement prediction, comparing its performance with traditional state-of-the-art models. These works collectively demonstrate the increasing sophistication of models that integrate text and financial data, highlighting the potential for improving trading scenarios.
@@ 4 Datasets
4.1 Stastics
@@
News
Market Data
Alpha Factors
Fundamental Data
4.2 Data Split
5 Methodology
@@
In TradExpert, all expert LLMs are built on the LLaMA-2-7B backbone LLM Touvron et al. (2023b) and are supervised and fine-tuned using the LoRA mechanism Hu et al. (2022). Before training and fine-tuning, we preprocess the raw datasets to construct prompts, instructions, and ground-truth responses for each LLM. An overall description of the preprocessed datasets is demonstrated in Table 1. The details will be introduced in the following.
5.1 News Analyst
The News Analyst LLM is designed to analyze texts of news articles to predict stock movements. The prompt and instruction for fine-tuning the LLM are shown in Figure 7. The outputs from the News Analyst LLM include not only a prediction of the stock movement but also a reasoning of how the news article relates to the predicted movement in order to employ a Chain-of-Thought (CoT) Wei et al. (2022) reasoning approach. The ground-truth reasonings are pre-generated by the OpenAI GPT-4 API using instructions and prompts that incorporate the actual stock movements and the texts of news articles.
@@ 5.1 News Analyst
5.2 Market Analyst
The Market Analyst LLM focuses on analyzing historical OHLCV (Open, High, Low, Close, Volume) data to predict stock movements. However, time series data is inherently continuous and lacks the discrete token structure that LLMs are designed to process. This misalignment poses a significant challenge in effectively utilizing LLMs on time series. To this end, we utilize a reprogramming mechanism Jin et al. (2024) to reprogram the input financial time series into text prototype representations.
Formally, let an OHLCV data instance be $\mathbf{X}^{(i)}\in\mathbb{R}^{N\times T}$ which consists of $N$ variables across $T$ time steps. $\mathbf{X}^{(i)}$ is first divided and embedded into a sequence of patch embeddings $\left\{\mathbf{X}_{P}^{(i)}\in\mathbb{R}^{N\times L_{P}\times d_{m}}\right\}$, where $L_{P}$ and $d_{m}$ are the number of patches and the patch embedding dimension respectively. The patches are then reprogrammed using a collection of text prototypes $\mathbf{E}^{\prime}\in\mathbb{R}^{V^{\prime}\times D}$, which is achieved by linearly probing the LLM’s pre-trained word embedding $\mathbf{E}\in\mathbb{R}^{V\times D}$, where $V$ and $V^{\prime}$ are the size of the vocabulary of the LLM and the text prototypes ( $V^{\prime}\ll V$), and $D$ is the embedding dimension. The reprogrammed patches are generated using a multi-head cross-attention mechanism: $\mathbf{Z}_{k}^{(i)}=\text{Softmax}\left(\frac{\mathbf{Q}_{k}^{(i)}\mathbf{K}_{k}^{(i)\top}}{\sqrt{d_{k}}}\right)\mathbf{V}_{k}^{(i)}$, where query $\mathbf{Q}_{k}^{(i)}=\mathbf{X}_{P}^{(i)}\mathbf{W}_{k}^{Q}$, key $\mathbf{K}_{k}^{(i)}=\mathbf{E}^{\prime}\mathbf{W}_{k}^{K}$, and value $\mathbf{V}_{k}^{(i)}=\mathbf{E}^{\prime}\mathbf{W}_{k}^{V}$ for each head $k$. The reprogrammed embeddings $\mathbf{O}^{(i)}$ are obtained by aggregating the outputs from each attention head and projecting them to the hidden dimensions of the backbone LLM. Finally, the reprogrammed embeddings are augmented with a language description of statistics extracted from TSFresh Christ et al. (2018), serving as prompts for the Market Analyst. An example of instruction and prompt is shown in Figure 4.
@@ 5.3 Alpha Expert
For each stock, we first calculate the values of all alpha factors based on OHLCV data and then derive a comprehensive score via a LightGBM-based model Ke et al. (2017). Subsequently, we select Top-K alphas that contribute most significantly to this comprehensive score. Descriptions of these Top-K alphas are retrieved from the database and, along with the calculated values, are included in the prompts and instructions for the Alpha Expert.
5.4 Fundamental Analyst
@@ 5.5 General Expert
In ranking mode, used for stock trading, the General Expert LLM functions as a comparator to establish the ranking ability. Specifically, given the summarized reports of two stocks, the General Expert LLM would determine which stock is likely to perform better in the future. To generate a Top-K ranking of stocks, we employ a relaxed comparison-based sorting similar to BubbleSort: We initially compare every pair of stocks and count the number of wins for each stock. Subsequently, we sort these counts to establish the rankings for stocks. Although algorithms like QuickSort and vanilla BubbleSort offer fewer comparisons for Top-K selection on average $\mathcal{O}(N\log N)$ and $\mathcal{O}(N\cdot K)$, we propose to use this relaxed comparison-based sorting alogrithm with $\mathcal{O}(N^{2})$ due to the non-transitive nature of LLM-based comparator Liu et al. (2024). Therefore, more comparisons tend to yield more accurate rankings in practice.
@@ 6.1 Datasets
  • Benchmark Datasets: We use publicly available benchmark datasets in stock movement prediction research including CIKM18 Wu et al. (2018), ACL18 Xu & Cohen (2018), and BigData22 Soun et al. (2022) datasets.
@@
6.2 Experimental Setup
In our experiments, the four expert LLMs and the General Expert LLM are bulit on the LLaMA-2-7B bakcbone model Touvron et al. (2023b) and are finetuned via LoRA Hu et al. (2022) mechanism.
@@ 6.2 Experimental Setup
6.3 Baselines
For stock movement prediction, the baselines include: (1) Hybrid Models: StockNet Xu & Cohen (2018), ALSTM-W Soun et al. (2022), ALSTM-D Soun et al. (2022), SLOT Soun et al. (2022). 2) Large Language Models: GPT-4 Achiam et al. (2023), Gemini Team et al. (2023), LLaMA2-70B Touvron et al. (2023b), LLaMA3-8B Dubey et al. (2024), FinMA-7B Xie et al. (2023), FinGPT-LlaMA2-7B Yang et al. (2023a), InternLM-7B Cai et al. (2024), Falcon-7B Almazrouei et al. (2023), Mixtral-7B Jiang et al. (2023).
For stock trading simulation, the baselines include: (1)Traditional Models: Random Forest Breiman (2001), Decision Tree Loh (2011), SVM Cortes & Vapnik (1995). (2) Deep Learning Models: A2C Mnih et al. (2016), PPO Schulman et al. (2017), SARL Ye et al. (2020), EIIE Jiang et al. (2017), and DeepTrader Wang et al. (2021). To reduce computational costs in backtesting, we evaluated all methods on datasets with stocks on the DOW 30 list, a subset of the S&P 500, with around 30 stocks.
@@ 6.3 Baselines
6.4 Results
6.4.1 Stock Movement Prediction
6.4.2 Stock Trading Simulation
7 Ablation Study
The Impacts of Experts
@@ 7 Ablation Study
8 Conclusion
In this study, we introduced TradExpert, a novel framework that harnesses the power of LLMs to enhance stock trading strategies. By integrating multiple specialized LLMs, each focused on distinct aspects of financial data, TradExpert provides a comprehensive and nuanced analysis that significantly outperforms traditional financial models in practice. Looking ahead, our goal is to explore how to employ TradExpert in the high-frequency trading scenario and extend its capabilities to encompass a wider range of global markets.